OpenStudy (anonymous):

A triangle has the following measurements. What is a possible length for the third side? AB = 4, BC = 17, CA = ? 18 5 10 12

OpenStudy (anonymous):

18

OpenStudy (anonymous):

can't it possible be 5 as well if 18 is possible ?

OpenStudy (anonymous):

im pretty sure @surdawi is correct

OpenStudy (anonymous):

no a^2+b^2 =c^2 4^2 + 17^2 = 305; c = 17.5 if we assume that c is the 17 b^2 = c^2 - a^2 b^2 = 17^2 - 4^2 = 273 b = 16.5